1. Harry'sHarry's Hair Systems issued $ 17 comma 000$17,000 worth of bonds at 105105. The proceeds for Harry'sHarry's Hair Systems are ​$nothingm. 2. Evelyn'sEvelyn's Egg Company issued $ 120 comma 000$120,000 worth of bonds at 9898. The proceeds for Evelyn'sEvelyn's Egg Company are ​$nothingm. 3. Curly'sCurly's Cutlery Company issued $ 100 comma 000$100,000 worth of bonds at 89.589.5. The proceeds for Curly'sCurly's Cutlery Company are ​$nothingm.
Answer #1

1) Proceeds from bonds = 17000*1.05 = 17850

2) Proceeds from bonds = 120000*.98 = 117600

3) Proceeds from bonds = 100000*89.5 = 89500